OpenStudy (anonymous):

Solve by the elimination method 0.05x+0.25y=66 0.15x=0.05y=72

OpenStudy (amistre64):

which = sign is a + ?

OpenStudy (saifoo.khan):

it must b -

OpenStudy (saifoo.khan):

x = 532.5 y = 157.5

OpenStudy (anonymous):

oops I apologize its 0.15x+0.05y=72

OpenStudy (saifoo.khan):

:o

OpenStudy (saifoo.khan):

x = 420 y= 180

OpenStudy (amistre64):

0.05x +0.25y = 66 ; multiply this by -3 0.15x +0.05y = 72 -.15x -0.75y = -198 0.15x +0.05y = 72 ; now add them together ---------------------- -.7 = -126 ; divide out the -.7 to get y = -126/-.7 = 180

OpenStudy (amistre64):

when you know what y equals, we can just plug it into one of the original equations and solve for x